PREX/CREX Nov 6, 2018 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  555 144 462
 Bluejeans link: https://bluejeans.com/555144462
 Offline room:   CEBAF Center F324-325

Agenda

  1. Upcoming November 11-12 collaboration meeting
    • Registration link (Please register if you haven't yet)
  2. analyzer/DAQ update (Bob/Ye/Cameron)
    • Corrected a non-standard wiring of BPMB and corrected some mapping issues in the Pan Control.db (notes)
    • Japan can make plots of the Counting House DAQ in standalone mode (jaPAN github discussion)
    • CC: seven out of 40 VQWK channels are not working. CC: the cavity, stripline BPMs are in good channels. BPM sign conventions are the same as the HRS convention
    • CG: Chris C will look at the bad VQWKs to see if they can be repaired. DM: will bring two from ISU to setup for test in the spring.
    • DM: Tyler and Jack found a part of the motion system from PREX1 and we need to find that as well.
    • ET deadtime testing is coming along, Carl Timmer suggests using the Java monitorGui to watch event flow-through (notes, more notes)
  3. main detector pmt linearity update: Devi ELOG Entry 3563
    • DeA: timing settings are ok. We see some different behavior between 120 and 240 flip rate.
    • DeA: (slide 6) the non linearity for 6 and 10 nA show a very narrow range in HV that is reasonable. KK: 6 nA is not terrible since we can go to 600 V and we should be fine. DeA: the error bars are statistical uncertainty (RMS/mean)
    • DeA: (slides 7-10) show similar features.
    • DeA: comparing 120 and 240 at 6 and 10 nA shows something odd. KK: worry that there is something wrong, because I can’t imagine what would produce such behavior. DM: will look at faster flip rates to see what is going to happen.
    • DeA: will look at 30 Hz flip rate as well.
    • KK: we really need to understand what is happening before we move forward. Would like to have the option to run at 240 for PREX. DM: will explore the timing a bit more and different flipping rates. Hope to brainstorm at the collaboration meeting.
  4. new SAM insertion tube length spec: Dustin ELOG Entry 3564, Cameron (corrected a 2.2 cm sign error, sorry!) ELOG Entry 3562
    • DM: we are keeping within our 1mm closest tolerance. We will most likely have a taper on the outside (second picture, more realistic). The light guide will not change,
    • DM: this is for the 3.5 cm shorter light guide.
    • KK: we should make an estimate for the distance between the quartz and the central axis so that we can make sure we match all the simulations.
  5. SEE BPM noise issues: Cameron ELOG Entry 3566, Tom Powers' references in Entry 3565
    • CC: at the end of September the Tritium experiment saw some noise in the BPMs. The harp scans do not show significant noise in the positions.
    • CC: this stopped after some issues were resolved in the injector (beam loss related). The noise in fourier space was removed after that.
    • CC: Tom Powers thinks that this could be related to current modulations at very high rate 10kHz.


Present

Cameron C, Dave G, KK, Eric K, Rich H, Ryan R, Amali P, Chandan G, Victoria O, Dustin M, Jinlong Z, Tao Y, David A, Jay B, Paul S, Seamus R, Silviu D, Devi A, Guido U,

Excused

Sangwha Park, Kent P, Bob M